OpenStudy (anonymous):

Prove: 4 13file:///C:/Users/Student/Desktop/pics%20for%20plato/2a4de8556f04bfa775832f871dfdea80

OpenStudy (sleepyhead314):

@TheBleep Hey there and Welcome to OpenStudy! :) you can attach files with the blue "Attach File" button below :)

OpenStudy (anonymous):

and thxs :)

OpenStudy (sleepyhead314):

hmm I can't seem to open the file? :-/ if it is a picture, perhaps you can try ScreenShotting it instead :)

OpenStudy (anonymous):

Proof: Statement Reason 1. p || q given 2. 4 12 For parallel lines cut by a transversal, corresponding angles are congruent. 3. 4. 9 13 For parallel lines cut by a transversal, corresponding angles are congruent. 5. 4 13 Transitive Property of Congruence i have to find the missing step

OpenStudy (sleepyhead314):

if 4 = 12 _______ 9 = 13 then 4 = 13 they are combining the two somehow if it is transitive property if a=b and b=c, then a=c the middle of that should be 12 = 9 (I think)

OpenStudy (anonymous):

well if its not the right answer which im pretty sure it is thx for ur help :)

OpenStudy (anonymous):

i got another question if you can help if you want :)

OpenStudy (sleepyhead314):

I can try :) it's a little tough without a picture xD

OpenStudy (anonymous):

yea xD

OpenStudy (anonymous):

What is the missing reason in the proof? segment addition Congruent Segments Theorem Transitive Property of Equality Subtraction Property of Equality

OpenStudy (anonymous):

maybe you can now xD im figuring this stuff out :)

OpenStudy (sleepyhead314):

what they did was substitute CF = EB into AE + EB = CF + FD to get AE + CF = CF + FD then - CF -CF subtracted CF from both sides ------------------------------- AE = FD ^^^ Subtraction Property of Equality would be my best guess

OpenStudy (anonymous):

thx :D

OpenStudy (sleepyhead314):

if you need any more help, just tag me when I'm online :) @sleepyhead314

OpenStudy (anonymous):

kk